Calculate the speed it takes a celestial particle to move from Jupiter to Pluto.  



1
Expert's answer
2021-01-12T12:12:30-0500

The distance from the Sun to the Jupiter: 5.2 AU

The distance from the Sun to the Pluto: 39.5 AU

If it takes 2 days for a celestial particle to cover the distance between Jupiter and Pluto, its speed is


"v=\frac{d}{t}=\frac{d_P-d_J}{t},\\\space\\
v=\frac{(39.5-5.2)\cdot149.6\cdot10^6}{2\cdot24}=106.9\cdot10^6\text{ km/h}."